OpenStudy (anonymous):

The area of the front face of the analog clock shown below is 153.86 square inches. The length of the minute hand is 0.2 inches less than the radius of the front face. What is the length of the arc, rounded to the nearest hundredth of an inch, the minute hand makes when it moves from the number 8 to the number 2 on the clock?

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

Huh?

OpenStudy (jim766):

in this question we have 2 circles 1. the face of the clock 2. the circle the tip of the minute hand makes first we need to find the radis of the inner circle (the length of the minute hand) A = pi (r)^2 formula for area of a circle 153.86 = pi (r^2), now divide both sides by pi 153.86 = 3.14(r^2) ------ ------ 3.14 3.14 49 = r^2 7 = r this is the radis of the face of the clock. 7 - .2 = 6.8 is the length of the minute hand (the length of the min hand) you with me so far?

OpenStudy (anonymous):

Okay so the length of the arc is 6.8?

OpenStudy (anonymous):

Sorry mistaken thought area was the radius . Go with the explanation by @Jim766

OpenStudy (jim766):

no the radis of the inner circle is 6.8 now we need to find the circumference of this inner circle C = 2 (Pi) r C = 2(3.14) 6.8 what is the circumference of this circle?

OpenStudy (anonymous):

42.70, so then we divide by 2? and that gives me 21.35 as the length of the arc?

OpenStudy (jim766):

great!

OpenStudy (anonymous):

Thank you! I have one more question, can you help?

OpenStudy (jim766):

I can try...

OpenStudy (anonymous):

Jamie made a wax model of a rolling pin of diameter 8 cm. The rolling pin was shaped like a right circular cylinder with a right circular cone at each end as shown below. What was the total surface area of the rolling pin? Using complete sentences, describe the steps you used to calculate the surface area.

OpenStudy (jim766):

surface area of cylinder Surface Area = 2(pi r 2) + (2 pi r)* h surface area of cone SA = pi(r)s + pi(r^2)

OpenStudy (jim766):

in this type of problem, you find the area of the cylinder, then the 2 cones and add it together Start with the cylinder. We need radis and height can you get those from the diagram?

OpenStudy (anonymous):

Radius is 4, I'm not sure about height because of the two cones

OpenStudy (jim766):

if the cylinder was standing up, the height would be 12, do you see where I got that?

OpenStudy (anonymous):

Yes

OpenStudy (jim766):

SA = 2(3.14)(16) + 2 (3.14)(4 )(12) so what is the surface area of the cylinder?

OpenStudy (anonymous):

175.84

OpenStudy (jim766):

SA = 2(3.14)(16) + 2 (3.14)(4 )(12) Sa = 100.48 + 301.44 Sa = 401.92 please check me..

OpenStudy (anonymous):

Yeah sorry I missed the 4 in the above equation

OpenStudy (jim766):

now the cone SA = 3.14(4)(5)

OpenStudy (anonymous):

62.80

OpenStudy (jim766):

great, now there are 2 of them....

OpenStudy (jim766):

plus the area of the cylinder

OpenStudy (anonymous):

125.60 + 401.92 =527.52

OpenStudy (jim766):

nice job... make sense?

OpenStudy (anonymous):

Kind of, not really, lol it's starting to hurt my head. So the total surface area is 527.52 cm?

OpenStudy (jim766):

oh yeah, the units....are cm^2 527.52 cm^2 in area, units are always squared but that is it...

OpenStudy (anonymous):

Okay thank you!! :)

OpenStudy (jim766):

I know it is a little painful....keep working at it!